# Security Matrix for Multimodal Agents on Mobile Devices: A Systematic and Proof of Concept Study

### v1

@@ 1 Introduction
Existing studies mainly focus on agent systems in the Web scenario. For instance, Wu et al. [18] conducted adversarial attacks against MLLM-based Web agents targeting their image captioning models or the CLIP model [19]. Debenedetti et al. [20] created a dynamic environment for evaluating the security of LLM-based Web agent systems [20]. Unfortunately, the above attack techniques are not transferable to mobile agent systems. For instance, Wu et al. [18] added global adversarial perturbations into the input image to mislead the victim Web agent system, which is not practical for mobile agent system where the adversaries can hardly control all the pixels in the device screenshot.
@@ 2 Related Work
Despite the above progress, there is currently no study focusing on the security threats of the mobile MLLM agent systems. Existing security studies of MLLM agents mainly focus on Web scenarios. The work most similar to ours is that of Wu et al. [18], which conducts adversarial attacks to either mislead the MLLM-based Web agent workflow or degrade the agent system’s performance. Despite its success in black-box settings, it still has two limitations. First, the proposed attack assumes the victim system uses an image captioning model to enhance the understanding capability of MLLM, which may not be practical with the advancement of the reasoning capability of MLLM. Second, the attack perturbation is globally added into the screenshot image, which can hardly be implemented in some realistic application scenarios.

### v2

@@ 1 Introduction
Existing studies mainly focus on agent systems in the Web scenario. For instance, Wu et al. [18] conducted adversarial attacks against MLLM-based Web agents targeting their image captioning models or the CLIP model [19]. Debenedetti et al. [20] created a dynamic environment for evaluating the security of LLM-based Web agent systems [20]. Unfortunately, the above attack techniques are not transferable to mobile agent systems and how to achieve realistic attacks against mobile agent systems remains unexplored.
@@ 2 Related Work
Despite the above progress, there is currently no study focusing on the security threats of the mobile MLLM agent systems. Existing security studies of MLLM agents mainly focus on Web scenarios. The work most similar to ours is that of Wu et al. [18], which conducts adversarial attacks to either mislead the MLLM-based Web agent workflow or degrade the agent system’s performance. Despite its success in black-box settings, it cannot be directly utilized to attack mobile agent systems because of some restrictions. For instance, the proposed attack assumes the victim system uses an image captioning model to enhance the understanding capability of MLLM, which is not practical in the mobile agent domain.